OpenStudy (anonymous):

Give the least common multiple of 14 and 35. A. 7 B. 70 C. 350 D. 490

OpenStudy (anonymous):

14 = 7 * 2 35 = 7 * 5 LCM = 7*5*2 = 70

OpenStudy (anonymous):

 so clearly the answer is B